OpenStudy (anonymous):

If theta is in standard position and the terminal side of theta is in the quadrant III and on the line 4x-3y=0, find the exact value of sin theta.

OpenStudy (phi):

did you plot the line 4x-3y=0 ? First put it in slope-intercept form.

OpenStudy (anonymous):

no!

OpenStudy (anonymous):

the slope is 4/3

OpenStudy (phi):

slope intercept form is y= mx+b start with 4x-3y=0 add 3y to both sides to get 4x= 3y divide both sides by 3 (4/3) x = y or, y= (4/3) x can you plot this line?

OpenStudy (anonymous):

I got that! I have the graph drawn.

OpenStudy (phi):

 do you know the slope of the line?